OpenStudy (anonymous):

HELPPPP!!!! Use the rules of logarithms to expand log[(x+3)^4(x-5)^7]

Nnesha (nnesha):

alright familiar with the log rules ?

OpenStudy (anonymous):

not really. I'm new to them!

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to expan you can change division to subtraction product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition <----> multiplication power rule \[\large\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

expand*

OpenStudy (anonymous):

whoa. Okay.

OpenStudy (anonymous):

that's confusing. :(

Nnesha (nnesha):

so which rule would you apply first ?

Nnesha (nnesha):

\[\log[(x+3)^4 (x-5)^7] \] is same as \[\log[(x+3)^4* (x-5)^7] \]

OpenStudy (anonymous):

okay, which one is easier?

Nnesha (nnesha):

what do you mean ??

Nnesha (nnesha):

quotient rule\[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\] to expan you can change division to subtraction product rule \[\large\rm log_b x + \log_b y = \log_b( x \times y )\] addition <----> multiplication first you should apply one of these rule

Nnesha (nnesha):

there is a multiplication sign in the equation so would you apply product rule or quotient rule ?

OpenStudy (anonymous):

okay sorry how do i do that?

OpenStudy (anonymous):

i have no idea tbh

Nnesha (nnesha):

here is an example \[\log (x *2) \rightarrow \rm expand ~~~\log (x) + \log(2)\] it was multiplication to expand i changed it to addition

Nnesha (nnesha):

\[\log[(x+3)^4* (x-5)^7] \] how would you split it into two logs ?

OpenStudy (anonymous):

x+3^4 then x-5^7 ?

Nnesha (nnesha):

ye but what sign would be in between (x+3)^4 and (x-5)^7

Nnesha (nnesha):

\[\log (\color{ReD}{x} \color{green}{*}\color{blue}{2}) \rightarrow \rm expand ~~~\log (\color{ReD}{x}) \color{green}{+} \log(\color{blue}{2})\] mybe colors help u to understand these :=)

OpenStudy (anonymous):

*? or +?

OpenStudy (anonymous):

im sorry, im trying!! :(

Nnesha (nnesha):

\[\log (\color{ReD}{(x+3)^4} \color{green}{*}\color{blue}{(x-5)^7}) \]

Nnesha (nnesha):

*(multiplication sign like x )

Nnesha (nnesha):

since there is a multiplication sign you should convert it to addition (+ sign )

OpenStudy (anonymous):

so log(x+3)^4 + (x-5)^7 ? @Nnesha

Nnesha (nnesha):

no we should split it into two logs... \[\large\rm \color{reD}{ \log}(x+3)^4 +\color{Red}{log} (x-5)^7 \]

Nnesha (nnesha):

now apply the power rule

Nnesha (nnesha):

power rule \[\large\rm log_b x^\color{ReD}{y} = \color{reD}{y} \log_b x\]

OpenStudy (anonymous):

I'm seriously SO confused right now.. @Nnesha

Nnesha (nnesha):

what part you don't understand ???

Nnesha (nnesha):

here is an example \[\rm \log x^\color{Red}{2}\] would be \[\rm \color{Red}{2} logx\]

Nnesha (nnesha):

just move the exponent at front of the log

OpenStudy (anonymous):

so .. 4logx?

Nnesha (nnesha):

well the example i gave u there base is just `x` what's the base in your question ?

Nnesha (nnesha):

here is another example \[\large\rm \log (x)^\color{Red}{a} + \log (y)^\color{reD}{z} \rightarrow ~~~\color{red}{a}log(x) + \color{Red}{z}log(y)\]

OpenStudy (anonymous):

I don't understand any of this. i don't even know what to do.

Nnesha (nnesha):

there any 3 rules to expand log \[\large\rm log_b x - \log_b y = \log_b \frac{ x }{ y}\]\[\large\rm log_b x + \log_b y = \log_b( x \times y )\]\[\large\rm log_b x^y = y \log_b x\] that's it

Nnesha (nnesha):

look at 2nd rule \[\large\rm \color{Red}{\log}( x \times y )=\color{ReD}{log}~(x) + \color{red}{log}{(y)}\] in other words distribute parentheses by log and then just change the multiplication sign by + sign